OpenStudy (judson_rockets):

What is the volume of a cube, if the length of one side is 7 feet? A. 343 cubic feet B. 49 cubic feet C. 21 cubic feet D. 294 cubic feet

OpenStudy (abdullahm):

Volume of cube = \(\bf side \times side \times side = \boxed{\bf{side^3}}\)

OpenStudy (abdullahm):

And the length of your side is 7, so simply plug it in!

OpenStudy (judson_rockets):

21

OpenStudy (orange_):

343

OpenStudy (abdullahm):

@Judson_rockets That's wrong. What is 7 * 7 * 7 = ?? 7^3 is not the same thing as 7 * 3.

OpenStudy (judson_rockets):

343

OpenStudy (abdullahm):

And that is your final answer!

OpenStudy (judson_rockets):

Thank @Orange_ and @AbdullahM for helping

OpenStudy (abdullahm):

It was my pleasure!

OpenStudy (orange_):

Mine too! :D